# FPT approximations for Capacitated Sum of Radii and Diameters

### v1

@@ 2 Overview of Techniques
In this section, we highlight our conceptual and technical contributions. Due to space constraints, we begin with our main theorem (Theorem 3) and present the key ideas behind our algorithm for non-uniform capacities in Section 2.1. Subsequently, in Section 2.2, we outline the techniques used to establish our hardness result. Finally, in Section 2.3, we delve into the algorithmic ideas for uniform capacities. Note however, that from a pedagogical viewpoint, it is easier to begin reading first Section 2.3, followed by Section 2.1. We now set up basic notations required for the exposition.
@@ 2.1 Non-Uniform Capacitied Sum of Radii
Consider the following simple and natural strategy of processing clusters in $\mathcal{C}$ iteratively (see Algorithm 1 for pseudo-code). Initially all the clusters in $\mathcal{C}$ are unprocessed. We will process the clusters in $\mathcal{C}$ in non-decreasing order of their radii, denoted as $C_{1},\cdots,C_{k}$. Let $U_{1},\cdots,U_{k}$ be the corresponding cluster capacities. Consider the first iteration when we process $C_{1}$. Suppose we could find a dense ball $D_{1}={\sf ball}(y_{1},r_{1})$ of radius $r_{1}$ in $P$ such that $|D_{1}|\geq|C_{1}|$. Then, we could create a cluster $C^{\prime}_{1}\subseteq D_{1}$ of size $|C_{1}|$ [^10] Even though $|C_{1}|$ is unknown, in retrospect, we will be able to use Remark 1 to obtain such a clustering. with radius $r_{1}$, matching the optimal cost. To make this cluster permanent, we delete the points from $D_{1}$ to prevent them from being reassigned in later iterations. However, this could create problems — (i) $C^{\prime}_{1}$ may contain points from other clusters, so in the future iterations when we process these clusters then the densest ball may not have enough points, and (ii) we need to make sure the points of $C_{1}$ are taken care by some cluster in our solution. Our algorithm is based on the following two key ideas that handle these two issues:
@@
Case $1$: Hop through the anchors. Suppose we are lucky, and it turns out that $r_{j}\leq r_{i}$. Then, $y_{i}$ can serve all the points of $C_{i}$ within radius $3r_{i}+2r_{j}\leq 5r_{i}$. See Figure 1(b), where $C_{j}=C_{3}$. We now mark $C_{i}$ as processed. Note that, in this case, we do not delete any points.
@@
Case $2$: Otherwise “Exchange”. Suppose we are not lucky, and it turns out that $r_{j}>r_{i}$ (see fig. 1(c) where $r_{i}=r_{5}>r_{8}=r_{j}$). In this case, our basic approach is to open the cluster $C_{i}$ at $D_{i}$, and thus relieving the clusters intersecting $D_{i}$ from their responsibility for these points. In exchange, these clusters will become responsible for $C_{i}$ points. The crux is that as the anchor $C_{j}$ has the minimal radius among the clusters intersecting $D_{i}$, each such cluster $C_{q}$ has radius larger than $r_{i}$. Thus from $C_{q}$’s perspective, the points of $C_{i}$ are “nearby”. Hence $C_{q}$ can accept responsibility for a number of $C_{i}$ points proportional to the number of $C_{q}$ points taken (and thus everybody is taken care of).
@@
In more details: we create a new cluster $C^{\prime}_{i}\subseteq D_{i}$ out of the points in $D_{i}$ (recall $|D_{i}|\geq|C_{i}|$ and see footnote 10). To make this assignment permanent, we delete $D_{i}$, and mark $C_{i}$ as processed. However, we end up taking points from other clusters (and also end up deleting points from these clusters). The key observation is that as we process the clusters in non-decreasing order and $r_{j}>r_{i}$, $D_{i}$ does not contain points from clusters processed via Case $1$. Since we maintain the invariant that the points of unprocessed clusters are not deleted, we have to mark the clusters intersecting $D_{i}$ as processed. However, before marking these clusters as processed, we need to ensure that they are accounted for in our solution. To this end, we use a novel idea of exchanging points. Since we have created a new cluster $C^{\prime}_{i}$ out of $D_{i}$, any cluster $C_{j^{\prime}}$ that has lost, say, $n_{j^{\prime}}>0$ points in this process, can instead claim back $n_{j^{\prime}}$ points from the original cluster $C_{i}$ (depicted by islands in $C_{5}$ in Figure 1(c)) by paying slightly more cost since, in this case, we have $r_{i}<r_{j}\leq r_{j^{\prime}}$. Here, we use the fact that $C_{j}$ is the anchor for $D_{i}$, and hence $r_{j}$ is the smallest radius intersecting $D_{i}$. We call such clusters partitioned clusters. Specifically, the radius of the partitioned $C_{j^{\prime}}$ is at most $\delta(o_{j^{\prime}},y_{i})+\delta(y_{i},p)\leq(r_{j^{\prime}}+r_{i})+(r_{i}+2r_{j}+2r_{i})\leq 7r_{j^{\prime}}$, for $p\in C_{i}$, as $r_{j}\leq r_{j^{\prime}}$. See Figure 1(c) for an illustration. Note that a cluster $C_{j^{\prime}}$ can be partitioned multiple times by different $C_{i}$’s, as shown in Figure 1(d). However, $o_{j^{\prime}}$ can still serve all the points of the modified $C_{j^{\prime}}$ within cost $7r_{j^{\prime}}$, since, in this case, the radii of the clusters partitioning $C_{j^{\prime}}$ are strictly less than $r_{j^{\prime}}$. While it is hard to find $o_{j^{\prime}}$, in the technical Section 4, we show how to find another point that can serve all the points of the partitioned cluster $C_{j^{\prime}}$ within radius $7r_{j^{\prime}}$.
@@
Figure 1: Figure 1(a) shows a dense ball ${\sf ball}(y,r_{5})$, colored in black, that is not good for cluster $C_{5}$, because ${\sf ball}(y,r_{5}+2r_{8})$ is disjoint from $C_{5}$. Hence, it is safe to (temporary) delete ${\sf ball}(y,r_{5}+2r_{8})$ without affecting cluster $C_{5}$, and recursively search for good dense ball for $C_{5}$ in the remaining points. On the other hand, the ${\sf ball}(y,r_{5})$ of Figure 1(b) is good dense ball for $C_{5}$. Furthermore, note that in this case the radius of the anchor $r_{3}\leq r_{5}$, and hence $y$ can serve all the points of $C_{5}$ within distance $3r_{5}+2r_{3}\leq 5r_{5}$. In Figure 1(c), the radius of the anchor is $r_{8}>r_{5}$, and hence the previous argument fails. However, in this case, since ${\sf ball}(y_{,}r_{5})$ is a good dense ball for $C_{5}$, so we can open a new cluster $C^{\prime}_{5}\subseteq{\sf ball}(y,r_{5})$ of appropriate size, instead of serving the points of $C_{5}$. Note that, in this case, the new cluster $C^{\prime}_{5}$ ends up taking points from clusters $C_{8},C_{9}$ and $C_{10}$. However, these clusters can claim back the lost points from the original cluster $C_{5}$ (depicted by islands in $C_{5}$), which is not too far from them. In this case, we say that $C_{8},C_{9}$ and $C_{10}$ are partitioned by $C^{\prime}_{5}$. Figure 1(d) shows cluster $C_{8}$ partitioned by $C_{2},C_{4}$ and $C_{5}$, and it has to collect its lost points the respective clusters.
2.2 Hardness of FPT-Approximation for Uniform Capacities
Our hardness result is based on gap-ETH (gap-Exponential Time Hypothesis), introduced by [MR17, Din16]. On a high level, gap-ETH says that there exists an $\varepsilon>0$ such that given an instance $\phi$ of $3$-SAT on $N$ variables and $M$ clauses, no algorithm running in time $2^{o(N)}$ can distinguish if there is an assignment to the variables of $\phi$ that satisfy all the clauses in $\phi$, or if every assignment satisfies $<(1-\varepsilon)M$ clauses in $\phi$, even when $M=O(N)$. In fact, using a result from Feige [Fei98], we can assume that $\phi$ is regular — each variable appears in exactly $5$ clauses and no variable occurs more than once in a clause. This key property of $\phi$ allows us to create an instance of CapSoR with uniform capacities.
Reduction from $3$-SAT. We reduce $\phi$ to an instance $\mathcal{I}$ of uniform CapSoR, whose points are described by the vertices of a graph $H=(V,E)$ and the metric is the shortest path distances on $H$. See Figure 2 for an example. We begin by partitioning the variables arbitrarily into $k$ groups $G_{1},\cdots,G_{k}$, each consisting of roughly $N/k$ variables. For each group $G_{i}$, and each assignment to the variables of $G_{i}$: $\alpha|_{G_{i}}:G_{i}\rightarrow\{0,1\}$, we create a vertex in $V$, and we call them assignment vertices. Additionally, for each group $G_{i}$, add (roughly) $2^{N/k}$ auxiliary vertices. Let $V_{i}$ denote the set of vertices corresponding to $G_{i}$ (including auxiliaries). Note that $|V_{i}|=2^{O(N/k)}$. We add all possible edges between $V_{i}$ vertices. For each clause, we create a vertex in $V$ and call it clause vertex. Let $V_{c}$ be the set of vertices corresponding to the clauses. We add all possible edges between $V_{c}$ vertices. Finally, for every pair of assignment and clause vertices, we add an edge between them of the partial assignment corresponding to the assignment vertex satisfies the clause corresponding to the clause vertex. We set the uniform capacity $U=|V_{i}|+5N/k$.
Figure 2: Example of the reduction from $3$-SAT to uniform capacitated Sum of Radii with $k=3$. Here we have a formula $\phi$ with $10$ clauses over $N=6$ variables. Each variable appears in exactly $5$ clauses. We create a graph whose corresponding shortest path metric will be the Sum of Radii instance. The variables are partition into $k=3$ groups of size $2$ each. For each group $G_{i}$ we add a vertex for each of the $4$ possible assignments to the corresponding variables (surrounded by rectangles). We add also $4$ auxiliary vertices to each group (on the left). All possible edges are added between the vertices in each group, and their auxiliary vertices. In addition, we add a vertex corresponding to each clause (on the right), and all possible edges between those vertices. Finally, we add an edge between every (partial) assignment to the clauses it satisfies.
For correctness, suppose that there is a satisfying assignment $\alpha$ to $\phi$, we claim that the cost of the optimal solution to $\mathcal{I}$ is at most $k$. Indeed, for every $V_{i}$, we open a cluster at the unique vertex corresponding to a partial assignment agreeing with $\alpha$. Let $X$ be this set (of centers), and we define the clusters by assigning a vertex to the closet center in $X$. Since $\alpha$ satisfies all the clauses in $\phi$, we have that every clause vertex in $V_{c}$ is at a distance $1$ from $X$. Each center $x_{i}\in X\cap V_{i}$ will be responsible for the points in $V_{i}$, and in addition to at most $5N/k$ clause vertices (as each variable belong to at most $5$ clauses), and thus overall to at most $U$ points. Hence the radii of each cluster is at most $1$, and the total cost is $k$.
For the other direction, we show that if there is a feasible solution $X$ with cost at most $(1+\varepsilon/66)k$ to $\mathcal{I}$, then there is an assignment $\alpha_{X}$ that satisfies more than $(1-\varepsilon)M$ clauses, thus yielding a gap between both the cases. The crux of the argument in this case is that the auxiliary vertices force $X$ to have the following important properties: (1) $X$ does not have zero radius clusters. (2) There are not too many centers in $X$ that serve points at distance more than $2$ from them. These two properties together ensure that essentially we have a single center per set $V_{i}$ with radius $1$, and thus we can recover an assignment $\alpha_{X}$ satisfying more than $(1-\varepsilon)M$ clauses.
@@ 3.2 Norm objectives
[…]
@@ 3.3.2 $\varepsilon$-approximation of radii and diameters
[…]
@@
[…]
@@
We assume that our algorithms have a power to make guesses in constant time. In particular, it can guess an element correctly from with probability $p(k)$ in constant time, for some computable function $p$. It can be easily verified that an algorithm for a minimization (maximization) problem with such guesses running in $f(k)n^{O(1)}$ time can be transformed into a randomized algorithm without guesses that runs in time $h(k)n^{O(1)}$, for some $h$, by replacing each guess either with a procedure that enumerates all elements of the sample space when the size of the sample space is bounded by a function of $k$ or the corresponding sampling procedure, and returning a minimum (maximum) cost solution.

### v2

@@ 2 Overview of Techniques
In this section, we highlight our conceptual and technical contributions. Due to space constraints, we begin with our main theorem (Theorem 3) and present the key ideas behind our algorithm for non-uniform capacities in Section 2.1. Then, in Section 2.2, we delve into the algorithmic ideas for uniform capacities. Note however, that from a pedagogical viewpoint, it is easier to begin reading first Section 2.2, followed by Section 2.1. We now set up basic notations required for the exposition.
@@ 2.1 Non-Uniform Capacitied Sum of Radii
Consider the following simple and natural strategy of processing clusters in $\mathcal{C}$ iteratively (see Algorithm 1 for pseudo-code). Initially all the clusters in $\mathcal{C}$ are unprocessed (colored red). We will process the clusters in $\mathcal{C}$ in non-decreasing order of their radii, denoted as $C_{1},\cdots,C_{k}$. Let $U_{1},\cdots,U_{k}$ be the corresponding cluster capacities. Consider the first iteration when we process $C_{1}$. Suppose we could find a dense ball $D_{1}={\sf ball}(y_{1},r_{1})$ of radius $r_{1}$ in $P$ such that $|D_{1}|\geq|C_{1}|$. Then, we could create a cluster $C^{\prime}_{1}\subseteq D_{1}$ of size $|C_{1}|$ [^11] Even though $|C_{1}|$ is unknown, in retrospect, we will be able to use Remark 1 to obtain such a clustering. with radius $r_{1}$, matching the optimal cost. To make this cluster permanent, we delete the points from $D_{1}$ to prevent them from being reassigned in later iterations. However, this could create problems — (i) $C^{\prime}_{1}$ may contain points from other clusters, so in the future iterations when we process these clusters then the densest ball may not have enough points, and (ii) we need to make sure the points of $C_{1}$ are taken care by some cluster in our solution. Our algorithm is based on the following two key ideas that handle these two issues:
@@
Case $1$: Hop through the anchors. Suppose we are lucky, and it turns out that $r_{j}\leq r_{i}$ (Algorithm 2 of  Algorithm 2). Then, $y_{i}$ can serve all the points of $C_{i}$ within radius $3r_{i}+2r_{j}\leq 5r_{i}$. See Figure 1(b), where $C_{j}=C_{3}$. We now mark $C_{i}$ as processed (color black). Note that, in this case, we do not delete any points.
@@
Case $2$: Otherwise “Exchange”. Suppose we are not lucky, and it turns out that $r_{j}>r_{i}$ (see fig. 1(c) where $r_{i}=r_{5}>r_{8}=r_{j}$). In this case, our basic approach is to open the cluster $C_{i}$ at $D_{i}$, and thus relieving the clusters intersecting $D_{i}$ from their responsibility for these points (see Algorithm 4). In exchange, these clusters will become responsible for $C_{i}$ points. The crux of the argument is that as the anchor $C_{j}$ has the minimal radius among the clusters intersecting $D_{i}$, each such cluster $C_{q}$ has radius larger than $r_{i}$. Thus from $C_{q}$’s perspective, the points of $C_{i}$ are “nearby”. Hence $C_{q}$ can accept responsibility for a number of $C_{i}$ points proportional to the number of $C_{q}$ points taken (and thus everybody is taken care of).
@@
In more details: we create a new cluster $C^{\prime}_{i}\subseteq D_{i}$ out of the points in $D_{i}$ (recall $|D_{i}|\geq|C_{i}|$ and see footnote 11). To make this assignment permanent, we delete $D_{i}$, and mark $C_{i}$ as processed (color black). However, we end up taking points from other clusters (and also end up deleting points from these clusters). The key observation is that as we process the clusters in non-decreasing order and $r_{j}>r_{i}$, $D_{i}$ does not contain points from clusters processed via Case $1$. Since we maintain the invariant that the points of unprocessed clusters are not deleted, we have to mark the clusters intersecting $D_{i}$ as processed (color black). However, before marking these clusters as processed, we need to ensure that they are accounted for in our solution. To this end, we use a novel idea of exchanging points. Since we have created a new cluster $C^{\prime}_{i}$ out of $D_{i}$, any cluster $C_{j^{\prime}}$ that has lost, say, $n_{j^{\prime}}>0$ points in this process, can instead claim back $n_{j^{\prime}}$ points from the original cluster $C_{i}$ (depicted by islands in $C_{5}$ in Figure 1(c)) by paying slightly more cost since, in this case, we have $r_{i}<r_{j}\leq r_{j^{\prime}}$. Here, we use the fact that $C_{j}$ is the anchor for $D_{i}$, and hence $r_{j}$ is the smallest radius intersecting $D_{i}$. We call such clusters partitioned clusters. Specifically, the radius of the partitioned $C_{j^{\prime}}$ is at most $\delta(o_{j^{\prime}},y_{i})+\delta(y_{i},p)\leq(r_{j^{\prime}}+r_{i})+(r_{i}+2r_{j}+2r_{i})\leq 7r_{j^{\prime}}$, for $p\in C_{i}$, as $r_{j}\leq r_{j^{\prime}}$. See Figure 1(c) for an illustration. Note that a cluster $C_{j^{\prime}}$ can be partitioned multiple times by different $C_{i}$’s (we track the clusters that partitioned $C_{j^{\prime}}$ by ${\textsf{Partitions-of}}(C_{j^{\prime}})$), as shown in Figure 1(d). However, $o_{j^{\prime}}$ can still serve all the points of the modified $C_{j^{\prime}}$ within cost $7r_{j^{\prime}}$, since, in this case, the radii of the clusters partitioning $C_{j^{\prime}}$ are strictly less than $r_{j^{\prime}}$. While it is hard to find $o_{j^{\prime}}$, in Section 5, we show how to find another point that can serve all the points of the partitioned cluster $C_{j^{\prime}}$ within radius $7r_{j^{\prime}}$ (see for loop at Algorithm 1 of Algorithm 1).
@@ 2.2 Uniform Capacities
2.3 Hardness of FPT-Approximation for Uniform Capacities
Our hardness result is based on a polynomial time reduction from Max $k$-Coverage (Max $k$-Cov), where, given a universe $\mathcal{G}$ of elements, a positive integer $k$, and $k$ collections $\{\mathcal{C}_{1},\dots,\mathcal{C}_{k}\}$ of subsets of $G$, the task is to find $\{S_{1},\dots,S_{k}\}$ such that $S_{i}\in\mathcal{C}_{i},i\in[k]$ that maximizes the number of elements covered. It is known that Max $k$-Cov is NP-hard to approximate to a factor better than (11/e)(1-1/e), even when every set in the collection has precisely $|\mathcal{G}|/k$ elements. Under gap-ETH, there is no $\textsf{FPT}(k)$ algorithm for Max $k$-Cov that approximates to a factor better than (11/e)(1-1/e), even when every set in the collection has precisely $|\mathcal{G}|/k$ elements
On a high level, given an instance of Max $k$-Cov, we create the set-element incidence graph of the instance. We then add edges between the vertices corresponding to the sets of the same collection. Let $G=(V,E)$ be the resulting graph. The points of our CapSoR instance corresponds to the vertices in $V$, and the distance metric is given by the shortest path distance in $G$. Finally, we set the uniform capacity $U=|\mathcal{G}|/k+m$. It is easy to see that when there are $k$ sets $\{S_{1},\dots,S_{k}\}$ such that $S_{i}\in\mathcal{C}_{i},i\in[k]$ that cover $\mathcal{G}$, then we can find a solution for CapSoR with cost $k$, by selecting the points corresponding to the sets $\{S_{1},\dots,S_{k}\}$ as the centers. For the No case, observe that any solution to CapSoR must have clusters of size precisely $U$ due to our construction. This simple observation forces the centers to belong to different collections, yielding the result.
Figure 1: Figure 1(a) shows a dense ball ${\sf ball}(y,r_{5})$, colored in black, that is not good for cluster $C_{5}$, because ${\sf ball}(y,r_{5}+2r_{8})$ is disjoint from $C_{5}$. Hence, it is safe to (temporary) delete ${\sf ball}(y,r_{5}+2r_{8})$ without affecting cluster $C_{5}$, and recursively search for good dense ball for $C_{5}$ in the remaining points. On the other hand, the ${\sf ball}(y,r_{5})$ of Figure 1(b) is good dense ball for $C_{5}$. Furthermore, note that in this case the radius of the anchor $r_{3}\leq r_{5}$, and hence $y$ can serve all the points of $C_{5}$ within distance $3r_{5}+2r_{3}\leq 5r_{5}$. In Figure 1(c), the radius of the anchor is $r_{8}>r_{5}$, and hence the previous argument fails. However, in this case, since ${\sf ball}(y_{,}r_{5})$ is a good dense ball for $C_{5}$, so we can open a new cluster $C^{\prime}_{5}\subseteq{\sf ball}(y,r_{5})$ of appropriate size, instead of serving the points of $C_{5}$. Note that, in this case, the new cluster $C^{\prime}_{5}$ ends up taking points from clusters $C_{8},C_{9}$ and $C_{10}$. However, these clusters can claim back the lost points from the original cluster $C_{5}$ (depicted by islands in $C_{5}$), which is not too far from them. In this case, we say that $C_{8},C_{9}$ and $C_{10}$ are partitioned by $C^{\prime}_{5}$. Figure 1(d) shows cluster $C_{8}$ partitioned by $C_{2},C_{4}$ and $C_{5}$, and it has to collect its lost points from the respective clusters.
3 Further Related Work
Capacitated Max-$k$-Diameter is a special case of capacitated $\ell_{p}$ norm of diameters problem when $p=\infty$. Its uncapacitated version turns out to be computationally very challenging. While there is a simple $2$-approximation for Max-$k$-Diameter [Gon85], it is NP-hard to approximate the problem better than factor $2$ in $\ell_{\infty}$ metric even for $k=3$ [Meg90]. Very recently, Fleischmann [FKK+23] showed that, even for $k=3$, the problem is NP-hard to approximate better than factors $1.5$ and $1.304$ when the points come from $\ell_{1}$ and $\ell_{2}$ metrics, respectively. Given that these hardness results hold for constant $k$ (i.e., even for $k=3$), they also translate to FPT lower bounds. Thus, these hardness results imply that no FPT algorithm can beat these lower bounds for Max-$k$-Diameter, unless $\textsf{P}=\textsf{NP}$. Of course, the corresponding capacitated version of Max-$k$-Diameter is a generalization of Max-$k$-Diameter, and hence these FPT-hardness of approximations also hold for the capacitated version. For upper bounds, it is interesting to note that the reduction from Sum of Radii to Sum of Diameters mentioned in footnote 10 works even when solving Sum of Radii w.r.t. soft assignments. Thus using [GJ23] one can obtain a $4$ FPT-approximation for the uniform Capacitated Max-$k$-Diameter problem, matching our Theorem 5 for this case.
@@ 4.2 Norm objectives
A norm $\|\cdot\|:\mathbb{R}^{k}\rightarrow\mathbb{R}_{\geq 0}$ is monotone if for every two vectors $(x_{1},\dots,x_{k}),(y_{1},\dots,y_{k})$ such that $\forall i,~|x_{i}|\leq|y_{i}|$, it holds that $\|(x_{1},\dots,x_{k})\|\leq\|(y_{1},\dots,y_{k})\|$. A norm $\|\cdot\|$ is symmetric if for every permutation $\pi:[k]\rightarrow[k]$ it holds that $\|(x_{1},\dots,x_{k})\|=\|(x_{\pi(1)},\dots,x_{\pi(k)})\|$. Interestingly, as pointed out by [CS19], a symmetric norm is not necessarily monotone. Indeed, consider the set $C\subseteq\mathbb{R}^{2}$, which is the convex hull of the points $\{(1,1),(-1,-1),(0,0.5),(0.5,0),(0,-0.5),(-0.5,0)\}$ (illustrated on the right), and define $\|x\|_{C}$ to be the smallest $\lambda$ such that $x/\lambda\in C$. Then $\|\cdot\|_{C}$ is a symmetric norm over $\mathbb{R}^{2}$, and $\|(0,0.5)\|_{C}=1$, while $\|(0.5,0.5)\|_{C}=0.5$, and thus $\|\cdot\|_{C}$ is not monotone.
@@ 4.3.2 $\varepsilon$-approximation of radii and diameters
In our algorithms, we will assume that we have a close approximations of the optimal radii/diameters. Consider a (unknown) multi-set $T^{*}=\{t^{*}_{1},\dots,t^{*}_{k}\}$ of $k$ non-negative reals such that $t^{*}_{1}\leq\dots\leq t^{*}_{k}$ and the largest entry $t^{*}_{k}$ is known. We say that a multi-set $T=\{t_{1},\cdots,t_{k}\}$ of non-negative reals is an $\varepsilon$-approximation of $T^{*}$ if for all $i\in[k]$, it holds that $t_{i}\geq t^{*}_{i}$, and $\|(t^{*}_{1},\dots,t^{*}_{k})\|\leq\|(t_{1},\dots,t_{k})\|\leq(1+\varepsilon)\|(t^{*}_{1},\dots,t^{*}_{k})\|$, where $\|\cdot\|$ is a monotone symmetric norm. The following lemma says that we can guess an $\varepsilon$-approximation in $2^{O(k\log(k/\varepsilon))}$ time, when the largest entry $t^{*}_{k}$ is known.
@@
Let $\delta=\varepsilon/k$. Given $t^{*}_{k}$, it is known that one compute (see [BLS23, JKY24]), in time $2^{O(k\log(k/\delta))}$, a list $\mathcal{L}$ of $k$-multi-set non-negative reals of size $2^{O(k\log(k/\delta))}$ such that there exists $(t_{1},\dots,t_{k})\in\mathcal{L}$ with $t^{*}_{i}\leq t_{i}\leq t^{*}_{i}+\delta\cdot t^{*}_{k}$, for $i\in[k]$. First, note that $\|(t^{*}_{k},\dots,t^{*}_{k})\|\leq\sum_{i=1}^{k}\|t^{*}_{k}\cdot e_{i}\|=k\cdot\|t^{*}_{k}\cdot e_{k}\|\leq k\cdot\|(t^{*}_{1},\dots,t^{*}_{k})\|$, where the first inequality follows from triangle inequality, the equality is because $\|\cdot\|$ is symmetric, and the last inequality follows due to monotonicity. Hence,
@@
We assume that our algorithms have a power to make guesses in constant time. In particular, it can guess an element correctly with probability $p(k)$ in constant time, for some computable function $p$. It can be easily verified that an algorithm for a minimization (maximization) problem with such guesses running in $f(k)n^{O(1)}$ time can be transformed into a randomized algorithm without guesses that runs in time $h(k)n^{O(1)}$, for some $h$, by replacing each guess either with a procedure that enumerates all elements of the sample space when the size of the sample space is bounded by a function of $k$ or the corresponding sampling procedure, and returning a minimum (maximum) cost solution.